### v1

@@ 3.2 The trajectory length $T$
The second theme in the Markov chain mixture modeling literature is the close link between the classification accuracy and the trajectory length $T$. Put simply, classifying short trajectories with the correct labels is hard. For example, Ramoni and coauthors [30, Sec. 5.3.1] describe numerical experiments with increasing trajectory lengths that lead to steadily increasing accuracy.
Researchers have tackled the issue of trajectory length by focusing on the extremes of short trajectories ($T=3$, [37, 38]) or long trajectories ($T\rightarrow\infty$, [41, 42, 36]). As the result of these investigations, there are now structured initializations for EM that use short trajectories of length $T=3$ [37, 38], but these approaches require massive amounts of data ($N=10^{8}$ trajectories to achieve classification error $<10\%$ [37]). At the other extreme, there are initialization algorithms for EM that ensure perfect parameter identification in the limit $T\rightarrow\infty$ [41, 36].
Sec. 5 provides further insight into the intermediate regimes where $T$ is neither short nor long. There is a nonasymptotic bound on the classification error which is valid for any Markov chain mixture model with any trajectory length $T$. The new bound indicates the importance of long trajectories, since the trajectory length leads to greater separation in Kullback-Leibler divergence and exponentially higher classification accuracy.
4 Variational EM algorithm for Markov chain mixtures
This paper proposes an extension of the variational EM algorithm, which can be used to model any time-series data set as a mixture of finite-state Markov chains. The current section describes the new method in detail and describes the steps needed for any user to apply the method to data.
As the first conceptual step, the user must reduce the data to a finite state space $\mathcal{Y}=\{1,\ldots,s\}$. The states can be chosen based on domain-area expertise or based on spectral clustering (Alg. 1), as is common in the Markov state modeling literature. The typical number of states is $s=10^{0}$$10^{2}$.
@@ 4 Variational EM algorithm for Markov chain mixtures
This model is newly presented here, but it combines standard Bayesian formulations for a Markov state model [43] and a general mixture model [44]. The model places uninformative $\operatorname{Dir}(\bm{1}_{s})$ prior distributions on the parameters $\bm{\nu}_{i}$ and the rows of the matrix $\bm{P}_{i}(\alpha,\cdot)$. The model places a different $\operatorname{Dir}(\bm{1}_{k}/k)$ prior distribution on the mixture parameters $\bm{\mu}$. This prior distribution is theoretically and empirically motivated, as it leads to a posterior distribution that automatically prunes extraneous mixture components by setting the corresponding mixture probabilities to $\mu(i)\approx 0$ [45, 19, 46].
Given the model (2) and the observed trajectories, the Bayesian posterior distribution is hard to compute exactly. However, the posterior can be efficiently approximated by a factorized distribution (see [12, Sec. 10.2]):
($\mathbb{P}(\{Z^{n},\bm{\mu},\bm{\nu}_{i},\bm{P}_{i}\})\approx Q_{1}(\{Z^{n}\})\cdot Q_{2}(\{\bm{\mu},\bm{\nu}_{i},\bm{P}_{i}\}).$
This factorization assumes the latent variables $\{Z^{n}\}$ are independent from the mixture parameters $\{\bm{\mu},\bm{\nu}_{i},\bm{P}_{i}\}$. Given this factorization, the best approximation that minimizes the Kullback-Leibler divergence with respect to the posterior is a sequence of mutually independent random variables:
$\bm{Z}^{n}$ $\sim\operatorname{Mult}(\bm{\gamma}^{n}),$ (3)
$\bm{\mu}$ $\sim\operatorname{Dir}(\bm{N}),$
@@
These random variables can be calculated using the variational EM algorithm in Alg. 3.
@@
$V^{n}(\alpha,\beta)$ $=\sum\nolimits_{0\leq t\leq T-1}\mathds{1}\{Y_{t}^{n}=\alpha,\,Y_{t+1}^{n}=\beta\}.$
@@
$\hat{N}(i)$ $=\frac{1}{k}+\sum\nolimits_{n=1}^{N}\hat{\mathbb{P}}\{Z^{n}=i\},$
$\hat{N}_{i}(\alpha)$ $=1+\sum\nolimits_{n=1}^{N}\hat{\mathbb{P}}\{Z^{n}=i\}\,U^{n}(\alpha),$
$\hat{N}_{i,\alpha}(\beta)$ $=1+\sum\nolimits_{n=1}^{N}\,\hat{\mathbb{P}}\{Z^{n}=i\}V^{n}(\alpha,\beta).$
@@
$\log\hat{\mu}(i)$ $=\psi(\hat{N}(i))-\psi\Bigl{(}\sum\nolimits_{j=1}^{k}\hat{N}(j)\Bigr{)},$
$\log\hat{\nu}_{i}(\alpha)$ $=\psi(\hat{N}_{i}(\alpha))-\psi\Bigl{(}\sum\nolimits_{\beta=1}^{s}\hat{N}_{i}(\beta)\Bigr{)},$
$\log\hat{P}_{i}(\alpha,\beta)$ $=\psi(\hat{N}_{i,\alpha}(\beta))-\psi\Bigl{(}\sum\nolimits_{\gamma=1}^{s}\hat{N}_{i,\alpha}(\gamma)\Bigr{)}.$
@@
7:     Update probabilities ^$\hat{\mathbb{P}}\{Z^{n}=i\}$ via
$\frac{\hat{\mu}(i)}{C_{n}}\prod\nolimits_{\alpha=1}^{s}\hat{\nu}_{i}(\alpha)^{U^{n}(\alpha)}\prod\nolimits_{\alpha,\beta=1}^{s}\hat{P}_{i}(\alpha,\beta)^{V^{n}(\alpha,\beta)},$
and choose $C_{n}$ so that $\sum\nolimits_{1\leq i\leq k}\hat{\mathbb{P}}\{Z^{n}=i\}=1$.
@@
$+\sum_{i=1}^{k}\log\biggl{[}\frac{B(\bm{1}_{s})}{B(\hat{\bm{N}}_{i})}\prod_{\alpha=1}^{s}\hat{p}_{i}(\alpha)^{\hat{N}_{i}(\alpha)-1}\biggr{]}$
$+\sum_{i=1}^{k}\sum_{\alpha=1}^{s}\log\biggl{[}\frac{B(\bm{1}_{s})}{B(\hat{\bm{N}}_{i,\alpha})}\prod_{\beta=1}^{s}\hat{p}_{i,\alpha}(\beta)^{\hat{N}_{i,\alpha}(\beta)-1}\Biggr{]},$
where $B(\bm{N})=(\prod_{i}\Gamma(N_{i}))\,/\,\Gamma(\sum_{i}N_{i})$.
@@
11:Return $\hat{\bm{N}},\hat{\bm{N}}_{i},\hat{\bm{N}}_{i,\alpha},\hat{Z}^{n}=\operatorname*{argmax}\limits_{1\leq i\leq k}\hat{\mathbb{P}}\{Z^{n}=i\}$.
Examining Alg. 3, there are a couple of key changes from the standard EM algorithm (Alg. 2). First, the user receives information about the parameter $\{\bm{\mu},\bm{\nu}_{i},\bm{P}_{i}\}$ in the form of Dirichlet distributions, which encode uncertainty information about the parameters. The user can extract a point estimate for any parameter by using the mean formula for the Dirichlet distribution, e.g.,
@@
The user can also extract error bars by using the variance formula for the Dirichlet distribution, e.g.,
@@
As the second change, variational EM automatically determines the number of mixture components. Upon convergence, only a small number of mixture parameters satisfy $\hat{N}(i)\geq 1$. The optimization drives all the remaining mixture parameters to the lowest possible value $\hat{N}(i)=1/k$. These components are not represented by any labels $\hat{Z}^{n}=i$, so they are effectively removed from the model. The tendency of variational EM to identify a parsimonious mixture model is called automatic relevance determination. It was observed heuristically in the papers [14, 18] and later justified based on asymptotic arguments in [19].
To apply Alg. 3 in practice, the user needs some way to initialize the parameter estimates. Since there is limited theoretical understanding of the best way to initialize variational EM, this paper uses a random sampling technique. The algorithm is by default run with $10^{2}$ independent initializations based on randomly sampling parameters ^$\hat{\mathbb{P}}\{Z^{n}=i\}$ for $i=1,\ldots,k$ from the probability simplex. After running Alg. 3 with different initializations, the user selects the optimal parameters $\{\hat{\bm{N}},\hat{\bm{N}}_{i},\hat{\bm{N}}_{i,\alpha},\hat{Z}^{n}\}$ that maximize the likelihood bound $\mathcal{L}$.
5 Limits on Markov chain mixture modeling
Markov chain mixture modeling cannot succeed when the trajectories are too short. This section provides a theoretical explanation for the high misclassification rate by presenting and interpreting the following theorem.
Theorem 1 (Misclassification error bound).
Consider any statistical estimator $\hat{Z}$ for the label $Z$ in the Markov chain mixture model (1). The estimator can depend on the observed data $\bm{Y}$ and the model parameters but not on the unknown label $Z$. Then, the misclassification error is at least
In this expression, the Kullback-Leibler divergence is defined as
The summation runs over all possible trajectories of length $T$, and
$\mathbb{P}_{i}(\alpha_{0},\ldots,\alpha_{T})=\nu_{i}(\alpha_{0})\prod\nolimits_{t=0}^{T-1}P_{i}(\alpha_{t},\alpha_{t+1}).$
Proof.
Given the observed data $\bm{Y}$, the conditional distribution of the component label $Z$ is
$\mathbb{P}\{Z=i\,|\,\bm{Y}\}=\frac{\mu(i)\mathbb{P}_{i}(\bm{Y})}{\sum_{j=1}^{k}\mu(j)\mathbb{P}_{j}(\bm{Y})}.$ (5)
The optimal classification algorithm selects the labels according to
@@ 5 Limits on Markov chain mixture modeling
Next, use the inequality $21-\alpha\geq(1-\alpha^{2})/$ 2 to obtain the lower bound
@@
$\geq\frac{1}{2}-\frac{1}{2}\sum\nolimits_{i=1}^{k}\mathbb{P}\{Z=i\,|\,\bm{Y}\}^{2}$
$=\frac{1}{2}\sum\nolimits_{i=1}^{k}\mathbb{P}\{Z=i\,|\,\bm{Y}\}(1-\mathbb{P}\{Z=i\,|\,\bm{Y}\}).$
@@
$\mathbb{P}\{Z\neq\hat{Z}\}=\sum\nolimits_{\bm{Y}}\mathbb{P}\{Z\neq\hat{Z}\,|\,\bm{Y}\}$ (6)
$\geq\frac{1}{2}\sum\nolimits_{i,\bm{Y}}\mathbb{P}\{Z=i,\bm{Y}^{n}\}(1-\mathbb{P}\{Z=i\,|\,\bm{Y}\})$
$=\frac{1}{2}\sum\nolimits_{i=1}^{k}\mu(i)\,\mathbb{E}\bigl{[}1-\mathbb{P}\{Z=i\,|\,\bm{Y}\}\,|\,Z=i\bigr{]}.$
The expression above turns out to be analytically tractable and is the key to the analysis.
@@
[…]
@@
$\mathbb{E}\biggl{[}\frac{\mathbb{P}_{j}(\bm{Y})}{\mathbb{P}_{i}(\bm{Y})}\,\bigg{|}\,Z=i\biggr{]}$ $=\sum\nolimits_{\bm{Y}}\mathbb{P}_{i}(\bm{Y})\frac{\mathbb{P}_{j}(\bm{Y})}{\mathbb{P}_{i}(\bm{Y})}$
$=\sum\nolimits_{\bm{Y}}\mathbb{P}_{j}(\bm{Y})=1.$
@@
Combine with eq. (6) to complete the proof. ∎
Thm. 1 determines the theoretical limits on classification accuracy in terms of the Kullback-Leibler divergence. The theorem extends to general mixture models, and it is apparently new in the statistics literature.
To interpret the theorem, it is instructive to consider what happens to the KL divergence as the trajectory length $T$ approaches infinity. If the Markov chain $i$ has a unique stationary measure $\bm{\pi}_{i}\in[0,1]^{s}$, then in the limit $T\rightarrow\infty$ the ergodic theorem for Markov chains [47, Sec. 1.10] guarantees
@@
$\rightarrow\sum\nolimits_{\alpha,\beta=1}^{s}\pi(\alpha)p_{i}(\alpha,\beta)\log\biggl{(}\frac{p_{i}(\alpha,\beta)}{p_{j}(\alpha,\beta)}\biggr{)}$
$=\sum\nolimits_{\alpha=1}^{s}\pi(\alpha)\operatorname{D}_{\rm KL}(p_{i}(\alpha,\cdot)\,\|\,p_{j}(\alpha,\cdot)).$
By this derivation, the KL divergence $\operatorname{D}_{\rm KL}(\mathbb{P}_{i}\,\|\,\mathbb{P}_{j})$ is approximately $T$ times the averaged Kullback-Leibler divergence $\operatorname{D}_{\rm KL}(p_{i}(\alpha,\cdot)\,\|\,p_{j}(\alpha,\cdot))$ over all the states. The linear dependence of the Kullback-Leibler divergence on $T$ motivates the following rule of thumb. If the trajectory length doubles, the KL divergence doubles, and the misclassification rate squares. For example, doubling the length of a trajectory can reduce the misclassification rate from $1$-in-$10$ to just $1$-in-$100$.
Figure 2: True parameters and parameter estimates for simulated trajectories with $k_{\mathrm{true}}=4$, $s=3$, $N=100$, and $T=30$. Top row shows true parameters. Bottom row shows parameter estimates from variational EM with a maximum of $k=10$ components. Even with limited data, VEM correctly identifies the number of components.

### v2

@@ 3.2 The Trajectory Length $T$
The second theme in the Markov chain mixture modeling literature is the close link between the classification accuracy and the trajectory length $T$. Put simply, classifying short trajectories with the correct labels is hard. For example, Ramoni and coauthors [13, Sec. 5.3.1] describe numerical experiments with increasing trajectory lengths that lead to steadily increasing accuracy.
Researchers have tackled the issue of trajectory length by focusing on the extremes of short trajectories ($T=3$, [35, 36]) or long trajectories ($T\rightarrow\infty$, [39, 40, 34]). A couple papers propose structured initialization methods for EM based on short trajectories of length $T=3$ [35, 36], but these approaches require massive amounts of data, for example, $N=10^{8}$ trajectories to achieve classification error $10\%$ [35]. At the other extreme, algorithms can ensure perfect parameter identification in the limit $T\rightarrow\infty$ [39, 34].
Sec. 4 provides further insight into the intermediate regimes where $T$ is neither short nor long, by bounding the classification error for any Markov chain mixture model with any trajectory length $T$. Since the KL divergence grows roughly linearly with $T$, the lower bound on the classification error decays roughly exponentially in $T$, indicating the importance of long trajectories.
3.3 Variational EM Algorithm for Markov Chain Mixtures
The current section describes the variational EM method for Markov chain mixtures in detail and outlines the steps for a user to apply the method to data.
As the first conceptual step, the user must reduce the data to a finite state space $\mathcal{Y}=\{1,\ldots,s\}$. The states can be chosen based on domain expertise or based on spectral clustering [28], as is common in the Markov state modeling literature. The typical number of states is $s=10^{0}$$10^{2}$.
@@ 3.3 Variational EM Algorithm for Markov Chain Mixtures
This Bayesian formulation combines standard Bayesian priors for Markov state modeling [41] and mixture modeling [42]. Specifically, the model places uninformative $\operatorname{Dir}(\bm{1}_{s})$ prior distributions on the parameters $\bm{\nu}_{i}$ and the rows of the matrix $\bm{P}_{i}(\alpha,\cdot)$, and it places a sparsity-promoting $\operatorname{Dir}(\bm{1}_{k}/k)$ prior distribution on the mixture parameter $\bm{\mu}$. The $\operatorname{Dir}(\bm{1}_{k}/k)$ prior leads to a posterior distribution that automatically prunes extraneous mixture components by setting the corresponding mixture probabilities to $\mu(i)\approx 0$ [43, 25, 44].
Given the model (2) and the observed trajectories, the Bayesian posterior distribution is hard to compute exactly. However, the posterior can be efficiently approximated by a factored distribution (see [21, Sec. 10.2]):
$\mathbb{P}(\{Z^{n},\bm{\mu},\bm{\nu}_{i},\bm{P}_{i}\})\approx Q_{1}(\{Z^{n}\})\,Q_{2}(\{\bm{\mu},\bm{\nu}_{i},\bm{P}_{i}\}).$
This factorization assumes the latent variables $\{Z^{n}\}$ are independent from the mixture parameters $\{\bm{\mu},\bm{\nu}_{i},\bm{P}_{i}\}$. Given this factorization, the approximation minimizing the Kullback-Leibler divergence with respect to the posterior is a sequence of mutually independent random variables
$\bm{\mu}$ $\sim\operatorname{Dir}(\bm{N}),$ (3)
@@
with parameters that can be optimized using the variational EM algorithm in Alg. 1.
@@
$V^{n}(\alpha,\beta)$ $=\sum_{t=0}^{T-1}\mathds{1}\{Y_{t}^{n}=\alpha,\,Y_{t+1}^{n}=\beta\}.$
@@
[…]
@@
$\log\hat{\mu}(i)$ $=\psi(\hat{N}(i))-\psi\Bigl(\sum_{j=1}^{k}\hat{N}(j)\Bigr),$
$\log\hat{\nu}_{i}(\alpha)$ $=\psi(\hat{N}_{i}(\alpha))-\psi\Bigl(\sum_{\beta=1}^{s}\hat{N}_{i}(\beta)\Bigr),$
$\log\hat{P}_{i}(\alpha,\beta)$ $=\psi(\hat{N}_{i,\alpha}(\beta))-\psi\Bigl(\sum_{\gamma=1}^{s}\hat{N}_{i,\alpha}(\gamma)\Bigr).$
@@
[…]
and choose $C_{n}$ so that $\sum_{1\leq i\leq k}\hat{\mathbb{P}}\{Z^{n}=i\}=1$.
@@
$+\sum_{i=1}^{k}\log\biggl[\frac{B(\bm{1}_{s})}{B(\hat{\bm{N}}_{i})}\prod_{\alpha=1}^{s}\hat{\nu}_{i}(\alpha)^{\hat{N}_{i}(\alpha)-1}\biggr]$
$+\sum_{i=1}^{k}\sum_{\alpha=1}^{s}\log\biggl[\frac{B(\bm{1}_{s})}{B(\hat{\bm{N}}_{i,\alpha})}\prod_{\beta=1}^{s}\hat{P}_{i}(\alpha,\beta)^{\hat{N}_{i,\alpha}(\beta)-1}\Biggr],$
where $B(\bm{N})=(\prod_{i}\Gamma(N_{i}))\,/\penalty 50\,\Gamma(\sum_{i}N_{i})$.
@@
[…]
Examining Alg. 1, there are a couple of key changes from the standard EM algorithm (Sec. 3.1). First, the parameters $\{\bm{\mu},\bm{\nu}_{i},\bm{P}_{i}\}$ are described by Dirichlet distributions, which encode point estimates and uncertainty intervals. The user can extract a point estimate for any parameter by using the expected value formula for the Dirichlet distribution, e.g.,
@@
The user can extract uncertainty intervals by using the variance formula for the Dirichlet distribution, e.g.,
@@
As the second change, variational EM automatically determines the number of mixture components. Upon convergence, only a small number of mixture parameters satisfy $\hat{N}(i)\geq 1$. The optimization drives all the remaining mixture parameters to the lowest possible value $\hat{N}(i)=1/k$. These components are not represented by any labels $\hat{Z}^{n}=i$, so they are effectively removed from the model. The tendency of variational EM to identify a parsimonious mixture model is called automatic relevance determination. It was observed in the papers [23, 24] and later justified based on asymptotic arguments in [25].
[…]
4 Theory: Limits on Markov Chain Mixture Modeling
Markov chain mixture modeling cannot succeed when the trajectories are too short. This section provides a theoretical explanation by proving and interpreting Theorem. 1.
Proof of Thm. 1.
Given the observed data $\bm{Y}$, the conditional distribution of the label $Z$ is
$\mathbb{P}\{Z=i\,|\,\bm{Y}\}=\frac{\mu(i)\mathbb{P}_{i}(\bm{Y})}{\sum_{j=1}^{k}\mu(j)\mathbb{P}_{j}(\bm{Y})}$ (5)
for each $i=1,\ldots,k$. The optimal classification algorithm selects the labels according to
@@ 4 Theory: Limits on Markov Chain Mixture Modeling
Next, use the inequality $1-\alpha\geq(1-\alpha^{2})/2$ to obtain the lower bound
@@
$\geq\frac{1}{2}-\frac{1}{2}\sum_{i=1}^{k}\mathbb{P}\{Z=i\,|\,\bm{Y}\}^{2}$
$=\frac{1}{2}\sum_{i=1}^{k}\mathbb{P}\{Z=i\,|\,\bm{Y}\}(1-\mathbb{P}\{Z=i\,|\,\bm{Y}\}).$
@@
$\mathbb{P}\{Z\neq\hat{Z}\}=\sum_{\bm{Y}}\mathbb{P}\{Z\neq\hat{Z}\,|\,\bm{Y}\}\,\mathbb{P}(\bm{Y})$
$\geq\frac{1}{2}\sum_{i,\bm{Y}}\mathbb{P}\{Z=i,\bm{Y}\}(1-\mathbb{P}\{Z=i\,|\,\bm{Y}\})$
$=\frac{1}{2}\sum_{i=1}^{k}\mu(i)\frac{\sum_{\bm{Y}}\mathbb{P}\{Z=i,\bm{Y}\}(1-\mathbb{P}\{Z=i\,|\,\bm{Y}\})}{\mu(i)}.$
The latter quantity is a conditional expectation
$\mathbb{P}\{Z\neq\hat{Z}\}$ (6)
$\geq\frac{1}{2}\sum_{i=1}^{k}\mu(i)\,\mathbb{E}\bigl[1-\mathbb{P}\{Z=i\,|\,\bm{Y}\}\,|\,Z=i\bigr].$
The conditional expectation turns out to be analytically tractable and allows us to derive the bound directly.
@@
$\begin{aligned} &=\mathbb{E}\biggl[\log\biggl(\frac{\mathbb{P}_{j}(\bm{Y})}{\mathbb{P}_{i}(\bm{Y})}\biggr)\,\bigg|\,Z=i\biggr]\\ &\quad-\mathbb{E}\biggl[\log\biggl(\frac{\mu(i)}{\mu(j)}+\frac{\mathbb{P}_{j}(\bm{Y})}{\mathbb{P}_{i}(\bm{Y})}\biggr)\,\bigg|\,Z=i\biggr].\end{aligned}$
@@
$\mathbb{E}\biggl[\frac{\mathbb{P}_{j}(\bm{Y})}{\mathbb{P}_{i}(\bm{Y})}\,\bigg|\,Z=i\biggr]$ $=\sum_{\bm{Y}}\mathbb{P}_{i}(\bm{Y})\frac{\mathbb{P}_{j}(\bm{Y})}{\mathbb{P}_{i}(\bm{Y})}$
$=\sum_{\bm{Y}}\mathbb{P}_{j}(\bm{Y})=1.$
@@
$\log\bigl(\mathbb{E}\bigl[1-\mathbb{P}\{Z=i\,|\,\bm{Y}\}\,|\,Z=i\bigr]\bigr)$
$\geq-\operatorname{D}_{\rm KL}(\mathbb{P}_{i}\,\|\,\mathbb{P}_{j})-\log\biggl(\frac{\mu(i)}{\mu(j)}+1\biggr)$
and therefore
@@
Combine with eq. (6) to conclude that the classification error is at least
This completes the proof. ∎
Two points are worth emphasizing. First, Theorem 1 is non-asymptotic: it applies to any finite trajectory length $T$ without requiring mixing or stationarity. It thus provides a baseline limit in the “many short trajectories” regime, isolating an intrinsic identifiability barrier.
Second, the $e^{-\operatorname{D}_{\rm KL}}$ scaling in the error bound reflects a classical connection between classification error and information divergences. The optimal error for distinguishing $\mathbb{P}_{i}$ from $\mathbb{P}_{j}$ is
$\sum_{\bm{Y}}\min\{\bm{\mathbb{P}}_{i}(\bm{Y}),\bm{\mathbb{P}}_{j}(\bm{Y})\}=1-d_{\rm TV}(\mathbb{P}_{i},\mathbb{P}_{j}),$
where the total variation distance is
$d_{\rm TV}(\mathbb{P}_{i},\mathbb{P}_{j})=\tfrac{1}{2}\sum_{\bm{Y}}|\mathbb{P}_{i}(\bm{Y})-\mathbb{P}_{j}(\bm{Y})|.$
The Bretagnolle–Huber inequality [26] then bounds
$1-d_{\rm TV}(\mathbb{P}_{i},\mathbb{P}_{j})\geq\tfrac{1}{2}e^{-\operatorname{D}_{\rm KL}(\mathbb{P}_{i}\,\|\,\mathbb{P}_{j})}$
Thm. 1 derives a comparable bound for the Markov mixture, making explicit how the KL divergence controls the overall classification error. This result complements recent work by Lee et al. [45], who prove high-probability error bounds and also find that KL divergence governs the clustering error.
To interpret the theorem, it is instructive to consider what happens to the KL divergence in the specific case of a Markov chain mixture as the trajectory length $T$ approaches infinity. If the Markov chain $i$ has a unique stationary measure $\bm{\pi}_{i}\in[0,1]^{s}$, then in the limit $T\rightarrow\infty$ the ergodic theorem for Markov chains [46, Sec. 1.10] guarantees
@@
$\rightarrow\sum_{\alpha,\beta=1}^{s}\pi_{i}(\alpha)P_{i}(\alpha,\beta)\log\biggl(\frac{P_{i}(\alpha,\beta)}{P_{j}(\alpha,\beta)}\biggr)$
$=\sum_{\alpha=1}^{s}\pi_{i}(\alpha)\operatorname{D}_{\rm KL}(P_{i}(\alpha,\cdot)\,\|\,P_{j}(\alpha,\cdot)).$
By this derivation, the KL divergence $\operatorname{D}_{\rm KL}(\mathbb{P}_{i}\,\|\,\mathbb{P}_{j})$ is approximately $T$ times the averaged Kullback-Leibler divergence $\operatorname{D}_{\rm KL}(P_{i}(\alpha,\cdot)\,\|\,P_{j}(\alpha,\cdot))$ over all the states. The linear dependence of the Kullback-Leibler divergence on $T$ has a practical implication: longer trajectories help substantially. Doubling $T$ can turn a hard problem into an easy one, for example reducing error from $10\%$ to $1\%$.
Figure 2: True parameters and parameter estimates for synthetic trajectories with $k_{\mathrm{true}}=4$, $s=3$, $N=100$, and $T=30$. Top row shows true parameters. Bottom row shows parameter estimates from variational EM with a maximum of $k=10$ components. Even with limited data, VEM correctly identifies the number of components.